OpenStudy (highschoolmom2010):

what values for, Θ(0≤Θ≤2π) satisfy the equation tan^2Θ=-3/2 sec Θ

OpenStudy (highschoolmom2010):

its not A

jhonyy9 (jhonyy9):

like a first step i think will be helpfully than you rewrite tan theta = sin theta / cos theta so sqauared what will be ?

OpenStudy (highschoolmom2010):

i really dont understand how to do this at all....

jhonyy9 (jhonyy9):

do you know sec theta what mean exactly ?

OpenStudy (templeguy):

Can you answer my question? I take a humor class because I wanna be a television writer. So they gave us this video to see but what is the joke from 12:34 - 12:50? http://www.dailymotion.com/video/x3hrv5o

jhonyy9 (jhonyy9):

have you learned about sec cosec ?

OpenStudy (highschoolmom2010):

no not really this was a random question plugged in our lesson and i just want to know how to do it

jhonyy9 (jhonyy9):

but really than you dont learned it again about these nothing so you dont can solving it never you need to know these above wrote what mean and so much again

jhonyy9 (jhonyy9):

tan^2 theta = sin^2 theta / cos^2 theta do you know this ?

OpenStudy (highschoolmom2010):

yes

jhonyy9 (jhonyy9):

looke this please http://www.math.com/tables/trig/identities.htm

jhonyy9 (jhonyy9):

moment please

jhonyy9 (jhonyy9):

@highschoolmom2010 please are you here ?

jhonyy9 (jhonyy9):

i believ that from above wrote website link you got it easy sec theta equal what ? yes ?

OpenStudy (sshayer):

\[\sec ^2 \theta=1+\tan ^2 \theta.\] make a quadratic in sec theta and solve it

jhonyy9 (jhonyy9):

sin^2 theta -3 ---------- = ---------- hope you understand how result this cos^2 theta 2 cos theta sin^2 theta = 1-cos^2 theta from formula sin^2 x +cos^2 x = 1 and you can divide both sides by 1/cos theta 1-cos^2 theta -3 ------------- = ----- cross multiplie cos theta 2 2(1-cos^2 theta) = -3costheta 2 -2cos^2 theta = -3cos theta sign cos theta = x 2 -2x^2 = -3x 2x^2 -3x -2 = 0 solve this quadratic for x D = 9 +16 =25 x_1,2 = (-2+/- 5)/4 = -7/4 and 3/4 @Hartmn your opinion about above wrote please ?

jhonyy9 (jhonyy9):

@hartnn please

OpenStudy (sshayer):

there is nothing wrong in it.

jhonyy9 (jhonyy9):

ty

OpenStudy (sshayer):

\[\sec ^2 \theta-1=-\frac{ 3 }{ 2 }\sec \theta\] \[2\sec ^2 \theta+3\sec \theta-2=0\] \[\frac{ 2 }{ \cos ^2 \theta }+\frac{ 3 }{ \cos \theta }-2=0\] or \[2+3\cos \theta-2\cos ^2\theta=0\] or\[2\cos ^2 \theta-3\cos \theta-2=0\] this is same as yours Mr.Johny

OpenStudy (sshayer):

\[2\cos ^2 \theta-4\cos \theta+\cos \theta-2=0\] \[2\cos \theta \left( \cos \theta-2 \right)+1\left( \cos \theta-2 \right)=0\] \[\left( \cos \theta-2 \right)\left( 2\cos \theta+1 \right)=0\] \[\cos \theta=2(rejected)~as~\left| \cos \theta \right|\le1\] \[\cos \theta=-\frac{ 1 }{ 2 }\] find theta

OpenStudy (sshayer):

\[\cos \theta=-\cos \frac{ \pi }{ 3 }=\cos \left( \pi-\frac{ \pi }{ 3 } \right),\cos \left( \pi+\frac{ \pi }{ 3 } \right)\] can you find theta now?

OpenStudy (michele_laino):

hint: using the definition of the function \(\tan \theta\) and \(\sec \theta\), we can rewrite the equation as follows: \[\begin{gathered} \frac{{{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} + \frac{3}{2}\frac{1}{{\cos \theta }} = 0 \hfill \\ \hfill \\ \frac{{2{{\left( {\sin \theta } \right)}^2} + 3\cos \theta }}{{2{{\left( {\cos \theta } \right)}^2}}} = 0 \hfill \\ \end{gathered} \] which is equivalent to these conditions: \[2{\left( {\sin \theta } \right)^2} + 3\cos \theta = 0,\quad \cos \theta \ne 0\] or these ones: \[2{\left( {\cos \theta } \right)^2} - 3\cos \theta - 2 = 0,\quad \cos \theta \ne 0\]